OpenStudy (anonymous):

solve for 0<θ<2(3.14[pi]): 9cos^((2))θ-2cosθ-3=0 give answer in terms of 3.14[pi]

OpenStudy (anonymous):

Treat \(\cos \theta\) as if it were a variable and apply the Quadratic formula.

OpenStudy (anonymous):

how so i have beeen struggling wit this question all night

OpenStudy (anonymous):

Just as normal.\[\cos \theta = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

and then wer do i go from here its a 6 points question

OpenStudy (anonymous):

What do you get as \(\cos \theta\)?

OpenStudy (anonymous):

cos theta=-29(-+)square root of 4+108